# A Simplified Proof for the Edge-Density of 4-Planar Graphs

### v1

@@ 1 Introduction
  • Case 2: One vertex-neighbor of $f$ is a 1-quadrilateral. W.l.o.g. let $f_{0}$ be a 1-quadrilateral.
    • Case 2.1: The vertex of $f_{0}$ is $v_{0}$ or $v_{1}$. By symmetry, assume that the vertex of $f_{0}$ is $v_{0}$. Observe that at most two vertex-neighbors of $f_{0}$ are possibly 0-pentagons (the third vertex-neighbor is the 1-triangle $t_{1}$). If $f_{0}$ does not contribute charge to both its wedge-neighbors, then it has an excess of $\frac{2}{9}$ after Step 3 and so $\operatorname{ch}_{4}(f)\geq 0$. Otherwise, we have the configuration depicted in Figure˜6(b) and $f$ is the only vertex-neighbor of $f_{0}$ that is a 0-pentagon. Here, $f_{0}$ does not contribute charge in Step 1 to its 1-neighbor that is not $t_{0}$ (if it is a 1-triangle, then another 2-triangle pays). Therefore, $\operatorname{ch}_{3}(f_{0})=\frac{4}{9}$ and $f_{0}$ can contribute its excess of $\frac{1}{9}$ charge to $f$ in Step 4.
    • Case 2.2: The vertex of $f_{0}$ is neither $v_{0}$ nor $v_{1}$. Observe that $f$ is the only vertex-neighbor of $f_{0}$ that is a 0-pentagon. As a second crossing of the wedges of $t_{0},t_{1}$ would allow a reduction to Case 1, we can assume that the 1-neighbors of $f_{0}$ are 2-triangles, see Figure˜6(c). Therefore, $f_{0}$ does not contribute charge in Step 1 and has an excess of at least $\frac{2}{9}$ after Step 3. Thus, $f$ is satisfied after Step 4.
  • Case 3: One vertex-neighbor of $f$ is a 0-quadrilateral. W.l.o.g. let $f_{0}$ be a 0-quadrilateral.
    • Case 3.1: One of $f_{1},f_{4}$ is a 0-quadrilateral. W.l.o.g. let $f_{1}$ be a 0-quadrilateral. Let $e$ be the edge crossing the wedge of $t_{1}$, see Figure˜7(a). As no homotopic multi-edges are allowed, w.l.o.g. the face $f^{\prime}$ that is a 1-neighbor of $t_{0}$ and a 0-neighbor of $f_{0}$ is a 2-face. But $f^{\prime}$ can not be a 2-triangle, which is a contradiction to Proposition˜2.8.
@@
    • Case 3.2: One of $f_{1},f_{4}$ is a 0-pentagon. W.l.o.g. let $f_{1}$ be a 0-pentagon. Observe that if $f_{1}$ has a vertex-neighbor $f^{\prime}\neq f$ that is a 0-pentagon, then $\operatorname{ch}_{3}(f_{1})\geq\frac{3}{9}$ and $f_{1}$ can contribute enough charge to $f$ in Step 4, see Figure˜7(b). Otherwise, $\operatorname{ch}_{3}(f_{1})=\frac{1}{9}$ implies $\operatorname{ch}_{4}(f)\geq 0$, and therefore we can assume for the critical case that all wedge-neighbors of $f_{1}$ are 1-triangles. Here, only the configuration depicted in Figure˜7(c) exists in that the wedge of $t_{1}$ is not crossed twice. In this configuration, $\operatorname{ch}_{3}(f_{1})=0$ and $\operatorname{ch}_{4}(f_{1})\geq\frac{1}{9}$, which is enough for $f$ in Step 5.
    • Case 3.3: Both of $f_{1},f_{4}$ are 1-triangles. Then $f_{2},f_{3}$ can not be 0-faces. Also, at most one of them is a 1-triangle, as otherwise there would be an edge homotopic to $e_{3}$, see Figure˜8(a). Therefore, $f_{2}$ or $f_{3}$ is a larger face (which must be a 1-quadrilateral by Propositions˜2.6, 2.8 and 2.10) and we can refer to Case 2.
@@
  • Case 4: One vertex-neighbor of $f$ is a 1-triangle. W.l.o.g. let $f_{0}$ be a 1-triangle so that the vertex of $f_{0}$ is $v_{0}$. Then $f_{4}$ is also a 1-triangle. Further, $f_{1}$ and $f_{3}$ are 0-pentagons, because otherwise they would be 1-triangles causing edges homotopic to $e_{1}$ and $e_{4}$. Finally, this implies that $f_{2}$ is a 0-pentagon, see Figure˜8(b). As $f_{2}$ does not contribute charge to $f_{1},f_{3}$ in Step 3, we have $\operatorname{ch}_{3}(f_{2})\geq\frac{3}{9}$. Here, at most three vertex-neighbors of $f_{2}$ are 0-pentagons and therefore $f$ can receive at least $\frac{1}{9}$ charge in Step 4.
@@
The combination of Proposition˜2.20 and Proposition˜2.22 closes the proof of Theorem˜1.1.

### v2

@@ 2 Definitions
  • Case 2: There is a vertex-neighbor of $f$ that is a 1-quadrilateral. W.l.o.g. let $f_{0}$ be a 1-quadrilateral.
    • Case 2.1: The vertex of $f_{0}$ is $v_{0}$ or $v_{1}$. By symmetry, assume that the vertex of $f_{0}$ is $v_{0}$. Observe that at most two vertex-neighbors of $f_{0}$ are possibly 0-pentagons (the third vertex-neighbor is the 1-triangle $t_{1}$). If $f_{0}$ does not contribute charge to both its wedge-neighbors, then it has an excess of $\frac{2}{9}$ after Step 3 and so $\operatorname{ch}_{4}(f)\geq 0$. Otherwise, we have the configuration depicted in Figure˜7(b) and $f$ is the only vertex-neighbor of $f_{0}$ that is a 0-pentagon. Here, $f_{0}$ does not contribute charge in Step 1 to its 1-neighbor that is not $t_{0}$ (if it is a 1-triangle, then another 2-triangle pays). Therefore, $\operatorname{ch}_{3}(f_{0})=\frac{4}{9}$ and $f_{0}$ can contribute its excess of $\frac{1}{9}$ charge to $f$ in Step 4.
    • Case 2.2: The vertex of $f_{0}$ is neither $v_{0}$ nor $v_{1}$. Observe that $f$ is the only vertex-neighbor of $f_{0}$ that is a 0-pentagon. As a second crossing of the wedges of $t_{0},t_{1}$ would allow a reduction to Case 1, we can assume that the 1-neighbors of $f_{0}$ are 2-triangles, see Figure˜7(c). Therefore, $f_{0}$ does not contribute charge in Step 1 and has an excess of at least $\frac{2}{9}$ after Step 3. Thus, $f$ is satisfied after Step 4.
  • Case 3: There is a vertex-neighbor of $f$ that is a 0-quadrilateral. W.l.o.g. let $f_{0}$ be a 0-quadrilateral.
    • Case 3.1: $f_{1}$ or $f_{4}$ is a 0-quadrilateral. W.l.o.g. let $f_{1}$ be a 0-quadrilateral. Let $e$ be the edge crossing the wedge of $t_{1}$, see Figure˜8(a). As no homotopic multi-edges are allowed, w.l.o.g. the face $f^{\prime}$ that is a 1-neighbor of $t_{0}$ and a 0-neighbor of $f_{0}$ is a 2-face. But $f^{\prime}$ can not be a 2-triangle, which is a contradiction to Proposition˜3.8.
@@
    • Case 3.2: $f_{1}$ or $f_{4}$ is a 0-pentagon. W.l.o.g. let $f_{1}$ be a 0-pentagon. Observe that if $f_{1}$ has a vertex-neighbor $f^{\prime}\neq f$ that is a 0-pentagon, then $\operatorname{ch}_{3}(f_{1})\geq\frac{3}{9}$ and $f_{1}$ can contribute enough charge to $f$ in Step 4, see Figure˜8(b). Otherwise, $\operatorname{ch}_{3}(f_{1})=\frac{1}{9}$ implies $\operatorname{ch}_{4}(f)\geq 0$, and therefore we can assume for the critical case that all wedge-neighbors of $f_{1}$ are 1-triangles. Here, only the configuration depicted in Figure˜8(c) exists in that the wedge of $t_{1}$ is not crossed twice. In this configuration, $\operatorname{ch}_{3}(f_{1})=0$ and $\operatorname{ch}_{4}(f_{1})\geq\frac{1}{9}$, which is enough for $f$ in Step 5.
    • Case 3.3: Both $f_{1}$ and $f_{4}$ are 1-triangles. Then $f_{2},f_{3}$ can not be 0-faces. Also, at most one of them is a 1-triangle, as otherwise there would be an edge homotopic to $e_{3}$, see Figure˜9(a). Therefore, $f_{2}$ or $f_{3}$ is a larger face (which must be a 1-quadrilateral by Propositions˜3.6, 3.8 and 3.10) and we can refer to Case 2.
@@
  • Case 4: There is a vertex-neighbor of $f$ that is a 1-triangle. W.l.o.g. let $f_{0}$ be a 1-triangle so that the vertex of $f_{0}$ is $v_{0}$. Then $f_{4}$ is also a 1-triangle. Further, $f_{1}$ and $f_{3}$ are 0-pentagons, as otherwise they would be 1-triangles causing edges homotopic to $e_{1}$ and $e_{4}$. Finally, this implies that $f_{2}$ is a 0-pentagon, see Figure˜9(b). As $f_{2}$ does not contribute charge to $f_{1},f_{3}$ in Step 3, we have $\operatorname{ch}_{3}(f_{2})\geq\frac{3}{9}$. Here, at most three vertex-neighbors of $f_{2}$ are 0-pentagons and therefore $f$ can receive at least $\frac{1}{9}$ charge in Step 4.
@@
The combination of Proposition˜3.20 and Proposition˜3.22 closes the proof of Theorem˜1.1.